Find the radius of the circle having centre at (2,1) whose one of the chord is a diameter of the circle x^2+y^2-2x-6y+6=0

Grade:12th pass

1 Answers

Arun
5 years ago
Dear student
 
 

x2+y2-2x-6y+6=0

or, (x2-2x+1)+(y2-6y+9)-4=0

or, (x-1)2+(y-3)2=22

thus the radius of the smaller given circle is 2 and its centre is (1,3).

distance b/w the two centres = (1+4)1/2 = 51/2

thus the radius of required circle = (5 + 4)1/2 = 3.